### v1

@@ Abstract
Existing regulatory frameworks for biomedical AI include robustness as a key component but lack detailed implementational guidance. The recent rise of biomedical foundation models creates new hurdles in testing and certification given their broad capabilities and susceptibility to complex distribution shifts. To balance test feasibility and effectiveness, we suggest a priority-based, task-oriented approach to tailor robustness evaluation objectives to a predefined specification. We urge concrete policies to adopt a granular categorization of robustness concepts in the specification. Our approach promotes the standardization of risk assessment and monitoring, which guides technical developments and mitigation efforts.
@@
The growing presence of biomedical foundation models (BFMs), including large language models (LLMs), vision-language models (VLMs), and others, trained using biomedical or de-identified healthcare data suggests they will eventually become integral to healthcare automation. Discussions on the risks of deploying algorithmic decision-making and generative AI in medicine have focused on bias and fairness. Robustness [1] is an equally important topic. It generally refers to the consistency of model prediction to distribution shift and is measured using either averaged or worst-case metrics. Robustness issues manifest in every stage of the model lifecycle (see Fig. 1a-b), leading to performance degradation, and more alarmingly, the generation of misleading or harmful content by imperfect users or bad actors [2]. The robustness of software also carries legal responsibilities [3], especially when it requires formal licensing or deployment or is classified as a medical device when used in high-stakes environments [4, 5].
@@
We examined over 50 existing BFMs (link to data) in preprints and publications covering different biomedical domains, about 32% of them contain no and about 32% use very limited robustness assessments (see also the repo). The most commonly presented evidence of model robustness is consistent performance across multiple datasets. Although a convenient proxy, consistent performance is not equivalent to a rigorous robustness guarantee because the relationships between datasets are generally unknown. To ensure constructive and beneficial use of BFMs in the future, we need to consider robustness evaluation in the model lifecycle and safety bottlenecks in the intended application settings [6]. Box 1 provides a structured list of robustness concepts that warrant consideration in biomedical domains. Inspired by test case prioritization in software engineering [7], we suggest designing effective robustness tests according to task-dependent robustness specifications constructed from priority scenarios (or priorities, see Fig. 1c) to facilitate test standardization while extending upon existing customized tests as building blocks. We introduce the background and motivation before our proposal using examples.
@@
Foundation model characteristics. The versatility of use cases and the exposure to complex distribution shifts are two major challenges of robustness evaluation (or testing) [8] for foundation models that differentiate from prior generations of predictive algorithms. The versatility comes from foundation models’ increased capabilities at inference time with knowledge injection through in-context learning, instruction following, the use of external tools (e.g. function calling) and data sources (e.g. retrieval augmentation), and with user steering of model behavior using specially designed prompts. These new learning paradigms blur the line between development and deployment stages and open up more avenues where models are exploited for their design imperfections.
@@
Average-case & Worst-case robustness define the robustness concepts based on either the averaged or worst-case performance metric of a model.
@@
Instance-wise robustness refers to the consistency of model performance to label-preserving changes at the level of a particular or the worst-case instance.
@@
Distribution shifts come from natural changes in the data distribution or intentional and sometimes malicious data manipulation (i.e. adversarial distribution shift) [8]. However, their distinction is becoming increasingly nuanced in the era of foundation models [9] due to the growing diversity of use cases. In the biomedical domains, natural distribution shifts manifest in changing disease symptomatology, divergent population structure, etc. Inadvertent text deletion or image cropping result in data manipulations, potentially leading to adversarial examples that alter model behavior. More elaborate attacks have been designed that involve targeted manipulation in model development and deployment [10, 11]. Poisoning attacks involve stealthy modification of training data, while in backdoor attacks, a specific token sequence (called trigger) is inserted during model training and activate during inference time [12]. Distribution shifts in the deployment stage result in the majority of failure modes, including input transforms applied to text (deletion, substitution, and addition including prompt injection, jailbreaks, etc) or images (noising, rotation, cropping, etc). Both natural distribution shifts and data manipulation yield out-of-distribution data [13] and they can have high domain-specificity or be created to target specific steps in the model lifecycle, creating complex deployment settings.
Robustness framework limitations. Aside from the challenges in scope, generating appropriate test examples for robustness evaluation are not often discussed. Two important robustness frameworks in ML, adversarial and interventional robustness, come from the security and causality viewpoints, respectively. The adversarial framework typically requires a guided search of test examples within a distance-bounded constraint such as the bounds by edit distance for text and Euclidean distance for image in Fig. 1c, yet there is no guarantee the test examples are sufficiently naturalistic to reflect reality. The interventional framework requires predefined interventions and a corresponding causal graph, which is often not immediately available for all tasks. Theoretical guarantees provided by these frameworks generally require justifications in the asymptotic limit and don’t necessarily translate into effective robustness in diverse yet highly contextualized deployment settings characteristic of the specialized domains [14, 15]. Because robustness testing (and hence certification) is critically dependent on the robustness framework of choice, we should design robustness tests more aligned with naturalistic settings that reflect the priorities in particular domains.
@@
Effective robustness evaluations requires a pragmatic robustness framework. We considers two aspects central to its specification: (i) the degradation mechanism behind the distribution shift, and (ii) the task performance metric which requires protection against the shift. The mechanistic understanding of a robustness failure mode requires establishing a connection between (i) and (ii), which is costly to account for every type of user interaction, or impractical when the users have insufficient information on development history or blackbox access. Moreover, multiple degradation mechanisms can simultaneously affect a particular downstream task.
@@
Technical robustness evaluations in ML have focused on tackling a simplified threat (threat-based) for obtaining certification, where a specific degradation mechanism guides the creation of test examples. Most adversarial and interventional robustness tests fit into this category [9], yet these approaches often target a considerably broader set of scenarios than meaningful in reality. From the efficiency perspective, taking a priority-based viewpoint [7] and focusing on retaining task performance under common degradation mechanisms anticipated in deployment settings is sufficient. Robustness tests based on simplified threat models and priorities are not mutually exclusive because accounting for realistic and meaningful perturbations (priority-based) overlaps with distance-bounded perturbations (threat-based), while the outcomes of priority-based tests should directly inform model quality. Fig. 1c contains two examples comparing threat- and priority-based robustness tests for text and image inputs, which illustrate the relationship between these two approaches for designing robustness tests.
We refer to the collection of priorities that demand testing for an individual task as a robustness specification. To contextualize robustness specification in naturalistic settings, we constructed two examples in Box 2 for an LLM-based pharmacy chatbot for over-the-counter (OTC) medicine and a VLM-based radiology report copilot for magnetic resonance imaging (MRI), both of which are attainable with existing research in BFM development. The specification contains a mixture of domain-specific (e.g. drug interaction, scanner information) and general aspects (e.g. paraphrasing, off-topic requests) that can induce model failures. The specification breaks down robustness evaluation into operationalizable units such that each can be converted into a small number of quantitative tests with guarantees. In reality, the test examples may come from augmenting or modifying the specified information in an existing data record [14, 15], such as a clinical vignette or case report. The specification can accommodate future capability expansion of models and updated risk assessments accordingly. We discuss below the feasibility of our proposal using existing and potential realizations of robustness tests of BFMs for major types of robustness in application settings (see Box 1).
@@
Box 2: Robustness specification Pharmacy chatbot for OTC medication The chatbot which can generate a medication request for patients has been extensively tested for robustness under the following scenarios: 1. Up to 12 turns of conversation on topics related to primary care medication 2. Knowledge of specialized vocabulary, including products of major pharmaceutical vendors 3. Partly missing patient information (may refuse the task if patient is unwilling to provide sufficient information) 4. Drug combinations, recommended intake quantities, and dosage limits 5. Adverse drug interactions and patient history 6. Stock availability and provider preference 7. Limited prescription authority (will refuse the task if asked for non-OTC or prescription drugs) 8. Paraphrasing and typo in instruction prompt 9. Off-topic requests and non-existent drugs MRI radiology report copilot The copilot that generates a report from one or a series of images has been extensively tested for robustness under the following scenarios: 1. Up to 15 turns of conversation including comparison of up to five images 2. Data from top-ten vendors of MRI scanners 3. Variability of MRI acquisition mode, image contrast, cropping or masking, and resolution 4. Common imaging artifacts (may refuse the task if image quality is too low) 5. Major human organs and body sections (may refuse the task if images given don’t resemble human anatomy) 6. Knowledge of specialized vocabulary 7. Spatial relations of human anatomy 8. Temporal ordering of and distinguish disease manifestation 9. Paraphrasing and typo in instruction prompt 10. Off-topic requests and non-MRI images
@@
Population structure. Explicit or implicit group structures are often present in biomedical and healthcare data, including prominent examples such as subpopulations organized by age group, ethnicity, or socioeconomic strata, medical study cohorts with specific phenotypic traits, etc. BFM-enabled cross-sectional or longitudinal studies for patient similarity analysis or health trajectory simulation may lead to group or longitudinal robustness issues when evaluating on incompatible populations. Group robustness assesses the model performance gap between the best- and the worst-performing groups, either identifiable through the label or hidden in the dataset. Testing group robustness may modify subpopulation labels in patient descriptions to gauge the change in model performance [20]. At a finer granularity, instance-wise robustness represents the performance gap between instances that are more prone to robustness failures than others, which are likely corner cases. It is important when the model deployment setting requires a minimum robustness threshold for every instance. Robustness testing in this context may use a balanced metric to reflect the impact of input modifications across individual instances.
Uncertainty awareness. The machine learning community typically distinguishes between aleatoric uncertainty, which comes from inherent data variability, and epistemic uncertainty, which arises from insufficient knowledge of the model in the specific problem context. Robustness tests against aleatoric uncertainty may assess the sensitivity of model output to prompt formatting and paraphrasing, while assessing robustness to epistemic uncertainty may use out-of-context examples [21] to examine if a model acknowledges the significant missing contextual information in domain-specific cases (e.g. providing the model with a chest X-ray and asking for a knee injury diagnosis). Additionally, uncertain information may also be directly verbalized in text prompts, a fitting scenario in biomedicine, to examine its influence on model behavior. Overall, the current generation of robustness evaluations haven’t yet included realistic uncertain scenarios often encountered in medical decision-making, although robustness against uncertainty is an important topic.
@@
Previous scenarios primarily consider assessing a monolithic model using single-criterion robustness tests. Specifying and testing robustness for more complex systems should also account for performance tradeoffs, model architecture, and user interactions. From the testing perspective, evaluating tradeoffs between various robustness criteria offers a more realistic view of a model’s aggregated robustness. Thus, multi-criteria robustness tests are essential to assess whether the model’s behavior reaches an optimal balance across these criteria. From the pipeline perspective, models integrated into a healthcare workflow can affect downstream biomedical outcomes. For example, using LLMs to summarize or VLMs to generate case reports may influence clinician decisions by emphasizing certain conditions or sentiments, affecting diagnoses or procedures. This highlights the need for considering robustness tests with humans in the loop and behavioral robustness in diverse interaction settings to assess the model’s impact on the care journey.
@@
From the model perspective, as modularity and maintainability become increasingly important, decision-making will be delegated to specialized subunits in a multi-expert (such as a mixture of experts) or multiagent system with a centralized coordinating unit. In these compound AI systems, each addressable subsystem is subject to testing and maintenance according to capability demand and regulatory compliance (see Fig. 1c). For example, Polaris [22] from Hippocratic AI features a multiagent medical foundation model which writes medical reports and notes as well as engages in low-risk patient interactions. Future systems with specialized units can mimic the group decision-making process in healthcare [23] to manage real-world complexities through enhanced reasoning and cooperative performance gain. Robustness tests for compound AI systems may consider different specifications for subsystems depending on the part-part and part-whole relationship in identifying bottlenecks and cascading effects associated with robustness failures.
Advancing concrete policies
to guide implementation
@@
Ensuring robustness for BFMs requires advancing regulatory policies for both AI and health information technology. Currently, the details for implementing robustness standards in AI regulations, such as the EU AI Act and the US Federal Artificial Intelligence Risk Management Act, are incomplete or insufficiently detailed. Existing health information technology regulations such as US-based HTI-1 final rule by Office of the National Coordinator focuses primarily on transparency and disclosures of the use of predictive decision support models yet lack detailed robustness requirements. The situation is in part due to the lack of a safety bare minimum for specific biomedical applications [5] and the evolving technological landscape, which can exacerbate the challenges laid out at the beginning of this Comment. Mandating a robustness specification proposed here provides a means to align policy objectives with practical implementation strategies, supporting evidence collection and enabling effective risk management throughout the model lifecycle. We advocate that robustness specifications (i) should seek community endorsement to gain a broad adoption; (ii) should consider the permissible tasks and user group characteristics due to the difference in user journeys; (iii) should inform regulatory standards such as in the construction of quantitative risk thresholds [24] or safety cases by enriching the failure mode taxonomy and improving their informativeness.
@@
Establishing robustness specifications with consensus from the research community will incentivize systematic evaluations by both model developers, research institutions, and independent third parties. These evaluations can inform proactive community engagement efforts, such as training end-users to recognize potential failures and apply mitigation strategies, which are critical to calibrating trust in human-AI interactions in high-stakes contexts. Concise, inclusive failure-reporting mechanisms should be established to allow users to provide timely feedback to the deployment team. Cross-referencing the specification with existing incident reporting mechanisms [6] will help pinpoint model vulnerabilities and guide targeted improvements or inform post-hoc adjustments to model behavior.

### v2

@@ Abstract
The rise of biomedical foundation models creates new hurdles in model testing and authorization given their broad capabilities and susceptibility to complex distribution shifts. We suggest tailoring robustness tests according to task-dependent priorities and propose to integrate granular notions of robustness in a predefined specification to guide implementation. Our approach facilitates the standardization of robustness assessments in the model lifecycle and connects abstract AI regulatory frameworks with concrete testing procedures.
Main text
@@
The growing presence of biomedical foundation models (BFMs), including large language models (LLMs), vision-language models (VLMs), and others, trained using biomedical or de-identified healthcare data, suggests they will eventually become integral to healthcare automation. Discussions on the risks of deploying algorithmic decision-making and generative AI in medicine have focused on bias and fairness. Robustness [1] is an equally important topic, which generally refers to the consistency of model prediction to distribution shifts. It is quantified using aggregated performance metrics, stratified comparisons across subsets of data, and worst-case performance. Robustness failures are an origin of the performance gap between model development and deployment, performance degradation over time, and more alarmingly, the generation of misleading or harmful content by imperfect users or bad actors [2]. The robustness of software also affects the legal responsibilities of providers [3] because the software may cause harm (e.g. misinformation, financial loss, or injury) to users or third parties or requires regulatory body authorization for deployment (e.g. medical devices) [4, 5].
Figure 1: Existing robustness tests used for biomedical foundation models. The treemap in a illustrates the topical areas of the BFMs looked at for this study. “General biomedical” indicates that the model is trained on general-purpose biomedical datasets and no domain specialization is emphasized in the model description. b shows the distributions of robustness tests (eval. = evaluation). Because multiple tests were conducted for some models, the total proportion in b is larger than unity.
@@
We examined over 50 existing BFMs covering different biomedical domains (see Fig. 1 and Supplementary Data 1). About 31.4% of them contain no robustness assessments at all. The most commonly presented evidence of model robustness is consistent performance across multiple datasets, which is adopted in 33.3% of the selected BFMs. Despite being a convenient proxy, consistent performance is not equivalent to a rigorous robustness guarantee because the relationships between datasets are generally unknown. Evaluations on shifted (5.9%) or synthetic data (3.9%), or data from external sites (9.8%) can be more effective but are not yet popular. To ensure the constructive and beneficial use of BFMs, we need to consider robustness evaluation across the model lifecycle and in intended application settings [6]. In biomedical domains, the various robustness concepts that warrant consideration (see Box 1 and the repo) motivate test customization. Inspired by test case prioritization in software engineering [7], which improves the cost-effectiveness of software testing by focusing on important test scenarios, we suggest designing effective robustness tests according to task-dependent robustness specifications constructed from priority scenarios (or priorities, see Fig. 2c) to facilitate test standardization while utilizing existing specialized tests as building blocks. Next, we introduce our proposal along with the background and motivation.
@@
Foundation model characteristics. The versatility of use cases and exposure to complex distribution shifts are two major challenges of robustness evaluation (or testing) [8] for foundation models that differentiate from prior generations of predictive algorithms. The versatility comes from foundation models’ increased capabilities at inference time with knowledge injection through in-context learning, instruction following, the use of external tools (e.g. function calling) and data sources (e.g. retrieval augmentation), and with user steering of model behavior using specially designed prompts. These new learning paradigms blur the line between development and deployment stages and open up more avenues where models are exploited for their design imperfections.
@@
Average-case & Worst-case robustness define the robustness concepts based on either the average or worst-case performance metric of a model.
@@
Instance robustness refers to the consistency of model performance to label-preserving changes at the level of a particular or the worst-case instance.
@@
Distribution shifts arise from natural changes in the data or intentional and sometimes malicious data manipulation (i.e. adversarial distribution shift) [8]. However, their distinction is increasingly nuanced in the era of foundation models [9] due to the growing number of use cases. Natural distribution shifts can manifest biomedically in changing disease symptomatology, divergent population structure, etc. Inadvertent text deletion or image cropping results in data manipulations, potentially leading to adversarial examples that alter model behavior.
More elaborate shifts have been designed by targeted manipulation in model development and deployment [10, 11] through the cybersecurity lens. Poisoning attacks involve stealthy modification of training data, while in backdoor attacks, a specific token sequence (called a trigger) is inserted during model training and activated during inference time [12]. Distribution shifts in the deployment stage result in the majority of failure modes, including input transforms applied to text (deletion, substitution, and addition, including prompt injection, jailbreaks, etc) or images (noising, rotation, cropping, etc). Both natural distribution shifts and data manipulation yield out-of-distribution data [13]. They can have high domain-specificity or be created to target specific aspects of the model lifecycle, resulting in complex origins that are hard to trace exactly.
Robustness framework limitations. Aside from the challenges in scope, how to generate appropriate test examples for robustness evaluation is not often discussed. Two important robustness frameworks in ML, adversarial and interventional robustness, come from the security and causality viewpoints, respectively. The adversarial framework typically requires a guided search of test examples within a distance-bounded constraint, such as the bounds established by edit distance for text and by Euclidean distance for image in Fig. 2c, yet there is no guarantee that the test examples are sufficiently naturalistic to reflect reality. The interventional framework requires predefined interventions and a corresponding causal graph, which is not immediately available for every task. Theoretical guarantees provided by these frameworks generally require justifications in the asymptotic limit and don’t necessarily translate into effective robustness in diverse yet highly contextualized deployment settings of specialized domains [14, 15]. Because robustness testing (and hence its associated guarantee) is critically dependent on the robustness framework of choice, we should design robustness tests that are more aligned with naturalistic settings and reflective of the priorities in corresponding domains.
@@
Effective robustness evaluations require a pragmatic framework. The two aspects central to its specification are: (i) the degradation mechanism behind a distribution shift, and (ii) the task performance metric that requires protection against the shift. Mechanistically understanding a robustness failure mode requires establishing a connection between (i) and (ii), which is costly when accounting for every type of user interaction or impractical when the users have insufficient information on model development history or blackbox access. Moreover, multiple degradation mechanisms can simultaneously affect a particular downstream task.
@@
Technical robustness evaluations in ML have generally been tackling simplified threats for obtaining statistical guarantees, where a specific degradation mechanism guides the creation of test examples. Most adversarial and interventional robustness tests fit into this category [9], often resulting in target a considerably broader set of scenarios than are meaningful in reality. From the efficiency perspective, taking a priority-based viewpoint [7] and focusing on retaining task performance under commonly anticipated degradation mechanisms in deployment settings is sufficient. Robustness tests based on simplified threat models and priorities are not mutually exclusive because accounting for realistic and meaningful perturbations (priority-based) has certain overlap with distance-bounded perturbations (threat-based), while the outcomes of priority-based tests should directly inform model quality. Fig. 2c contains two examples comparing threat- and priority-based robustness tests for text and image data inputs. It illustrates the relationship between these two approaches for designing robustness tests.
@@
Box 2: Robustness specification Pharmacy chatbot for OTC medication The chatbot which can generate a medication request for patients has been extensively tested for robustness under the following scenarios: 1. Up to 12 turns of conversation on topics related to primary care medication 2. Knowledge of specialized vocabulary, including products of major pharmaceutical vendors 3. Partly missing patient information (may refuse the task if patient is unwilling to provide sufficient information) 4. Drug combinations, recommended intake quantities, and dosage limits 5. Adverse drug interactions and patient history 6. Stock availability and provider preference 7. Limited prescription authority (will refuse the task if asked for non-OTC or prescription drugs) 8. Paraphrases and typos in an instruction prompt 9. Off-topic requests and non-existent drugs MRI radiology report copilot The copilot that generates a report from one or a series of images has been extensively tested for robustness under the following scenarios: 1. Up to 15 turns of conversation including comparison of up to five images 2. Data from top-four vendors of MRI scanners 3. Variability of MRI acquisition mode, image contrast, cropping or masking, and resolution 4. Common imaging artifacts (may refuse the task if image quality is too low) 5. Major human organs and body sections (may refuse the task if images given don’t resemble human anatomy) 6. Knowledge of specialized vocabulary 7. Spatial relations of human anatomy 8. Temporal ordering of distinguishing disease manifestations 9. Paraphrasing and typo in instruction prompt 10. Off-topic requests and non-MRI images
We refer to the collection of priorities that demand testing for an individual task as a robustness specification. To contextualize it in naturalistic settings, we constructed two examples in Box 2 for an LLM-based pharmacy chatbot for over-the-counter (OTC) medicine and a VLM-based radiology report copilot for magnetic resonance imaging (MRI), both of which are attainable with existing research in BFM development. The specification contains a mixture of domain-specific (e.g. drug interaction, scanner information) and general aspects (e.g. paraphrasing, off-topic requests) that can induce model failures. The specification breaks down robustness evaluation into operationalizable units such that each is convertible into a small number of quantitative tests with guarantees. In reality, the test examples may come from augmenting or modifying the specified information in an existing data record [14, 15], such as a clinical vignette or case report. The specification can accommodate the future capability expansion of models and risk assessment updates accordingly. We discuss below the feasibility of our proposal using existing and potential realizations of major types of robustness tests for BFMs in application settings (see Box 1).
@@
Population structure. Explicit or implicit group structures are often present in biomedical and healthcare data, including prominent examples such as subpopulations organized by age group, ethnicity, or socioeconomic strata, medical study cohorts with specific phenotypic traits, etc. BFM-enabled cross-sectional or longitudinal studies for patient similarity analysis or health trajectory simulation may lead to group or longitudinal robustness issues when evaluating on incompatible populations. Group robustness assesses the model performance gap between the best- and the worst-performing groups, either identifiable through the label or hidden in the dataset. Testing group robustness may modify subpopulation labels in patient descriptions to gauge the change in model performance [20]. At a finer granularity, instance robustness represents the performance gap between instances that are more prone to robustness failures than others, which are likely corner cases. It is important when the model deployment setting requires a minimum robustness threshold for every instance. Robustness testing in this context may use a balanced metric to reflect the impact of input modifications across individual instances.
Uncertainty awareness. The machine learning community typically distinguishes between aleatoric uncertainty, which comes from inherent data variability, and epistemic uncertainty, which arises from insufficient knowledge of the model in the specific problem context. Robustness tests against aleatoric uncertainty may assess the sensitivity of model output to prompt formatting and paraphrasing, while assessing robustness to epistemic uncertainty may use out-of-context examples [21] to examine if a model acknowledges the significant missing contextual information in domain-specific cases (e.g. presenting the model with a chest X-ray image and asking for a knee injury diagnosis). Additionally, uncertain information may also be directly verbalized in text prompts, a fitting scenario in biomedicine, to examine its influence on model behavior. Overall, the current generation of robustness evaluations haven’t yet included realistic uncertain scenarios often encountered in medical decision-making, although robustness against uncertainty is an important topic.
@@
Previous scenarios primarily consider assessing a monolithic model using single-criterion robustness tests. Specifying and testing robustness for more complex AI systems should also account for performance tradeoffs, model architecture, and user interactions.
Adaptation / Update Execution Cost Potential usage of robustness tests
Model weights Model provider high Helps choose and optimize the training method that yields models with a better robustness-accuracy tradeoff
Model architecture Model provider high Helps choose architectural parameters that are easier to attain a robust model than others when training on the same dataset
User prompt Model user low Helps select more robust prompt templates and improve the reliability of outputs
User confidence Model user medium Helps adjust and calibrate user confidence to deploy the model for appropriate tasks
Table 1: Robustness tests in the adaptation and update of BFM-based devices and services.
@@
Metrics and stakeholders. Evaluating tradeoffs between various robustness metrics and criteria offers a balanced view of a model’s robustness across different dimensions and through metric aggregation. These more comprehensive robustness tests are essential in assessing whether the model’s behavior reaches an optimal balance or is suitable for applications with distinct risk levels or stakeholders (see SI section 1). When models are integrated into a healthcare workflow, they can affect downstream biomedical outcomes. For example, using LLMs to summarize or VLMs to generate case reports may influence clinician decisions by emphasizing certain conditions or sentiments, affecting diagnoses or procedures. This highlights the need for considering robustness tests with the relevant stakeholder(s) in the loop and behavioral robustness across diverse interaction settings to assess the model’s impact on the care journey.
@@
Compound systems. As modularity and maintainability become increasingly important, decision-making will be delegated to specialized subunits in a multi-expert (such as a mixture of experts) or multiagent system [22] with a centralized coordinating unit. In these compound AI systems, each addressable subsystem is subject to testing and maintenance according to capability demand and regulatory compliance (see Fig. 2c). For example, Polaris [23] from Hippocratic AI features a multiagent medical foundation model that writes medical reports and notes as well as engages in low-risk patient interactions. Future systems with specialized units can mimic the group decision-making process in healthcare [24] to manage real-world complexities through enhanced reasoning and cooperative performance gain. Robustness tests for compound AI systems may consider different specifications for subsystems depending on the part-part and part-whole relationship in identifying bottlenecks and cascading effects associated with robustness failures.
Bridging policy with implementation
@@
Ensuring robustness for BFMs requires advancing regulatory policies for both AI and health information technology. Currently, the leading AI regulatory frameworks, such as the EU AI Act and the US Federal AI Risk Management Act, recognize the relation between natural and adversarial notions of robustness but contain insufficient details to guide implementation in domain-specific applications (see SI section 2). Existing health information technology regulations, such as US-based HTI-1 final rule by the Office of the National Coordinator, focus primarily on transparency and disclosures of the use of predictive decision support models, yet lack detailed robustness requirements. The situation is in part due to the lack of a safety bare minimum for specific biomedical applications [5] and the fast-evolving technological landscape, which can exacerbate the challenges laid out at the beginning of this Comment. These existing gaps make concrete community-endorsed standards on robustness even more important.
Considerations in implementation. Mandating robustness specifications according to the tasks and the biomedical domains (see Fig. 1-2) provides a means to map policy objectives onto real-world implementations. It also facilitates evidence collection and enables effective risk management throughout the model lifecycle. We advocate that robustness specifications (i) should seek community endorsement to gain a broad adoption; (ii) should consider the permissible tasks and user group characteristics due to the difference in user journeys; (iii) should inform regulatory standards such as in the construction of quantitative risk thresholds [25] or safety cases by enriching the failure mode taxonomy of BFMs and improving their informativeness. These considerations will facilitate the implementation of robustness specifications and ensure that their adoption is within shared interests.
@@
Community benefits. Establishing a consensus-driven robustness specification from the research community will incentivize systematic efforts by model developers, research institutions, and independent third parties. For model developers, robustness testing informs model selection and update. For the model deployment team and model users, robustness testing allows for identifying inference-time adjustments of prompt templates to improve the reliability of BFM applications. These potential uses of robustness tests are summarized in Table 1. Robustness specifications can also provide templates for failure-reporting procedures to allow users to provide timely feedback to the deployment team. Integrating robustness specifications with incident reporting mechanisms [6] facilitates the identification of model vulnerabilities and guides targeted improvements or informs post-hoc adjustments to model behavior. Their implementations can assist the training of end-users to recognize potential failures, calibrate user confidence, and enact mitigation strategies.